OpenStudy (anonymous):

N- 1/5=7/20 what is N? Is it 3/20, 11/20, 8/25,6/15?

OpenStudy (anonymous):

N - (1/5) = (7/20) N = (7/20) + (1/5) N = (7/20) + (4/40) N = ?

OpenStudy (anonymous):

ops, typo on the last one: N = (7/20) + (4/20) N = ?

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

N - (1/5) = (7/20) N = (7/20) + (1/5) N = (7/20) + (4/20) N = 11/20

OpenStudy (anonymous):

N=(1/5)+(7/20) N=(4+7)/20 11/20 sorry not 4/5

OpenStudy (anonymous):

Thanks mathmind, and nacylam! :D

OpenStudy (anonymous):

ops, 4N = 11/20 N = 44/20 N = 11/5

OpenStudy (anonymous):

u welcom

OpenStudy (anonymous):

damn... i keep massing it up.. nvm..

OpenStudy (anonymous):

Lol! well out of,, 3/20 ,,,, 11/20 ,,,, 8/25,,,, 6/15,,,, I thought it was 11/20

OpenStudy (anonymous):

that is happen to me to

OpenStudy (anonymous):

yea, that's the answer

OpenStudy (anonymous):

yes is 11/20

OpenStudy (anonymous):

Haha! Thanks can i ask you guys one more question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so 2/3 times N=12 so would it be 8,20,6,or 18? I thought 18

OpenStudy (anonymous):

yep 18

OpenStudy (anonymous):

(2/3)N = 12 N = 12* (3/2) N = 36 /2 = 18

OpenStudy (anonymous):

2N/3=12 2N=36 N=36/2 N=18

OpenStudy (anonymous):

YAY - feels special - Ty ty

OpenStudy (anonymous):

haha np

OpenStudy (anonymous):

XD im studying for a test so yeh !!!